d. 4 - Find the equation of the tangent line to the function () - 2 2x - I atx .
Answer #1

at 1=5 fo)= 2231-1 f'in)= 2x x2 = 2/27-1 Jzn-1 So Slope af the tangent lime at 91=5 is f5) fis)= 2. [t'a) is slope of fen] V2r5-1 Since the Slope of funtion fla) is f'ca) and the Value d.= feot6), the tangent line af an y function fe1) at a Paaticular valve af 1=H, is - fear) = f'cal(1-26) Here Ho =D5 So ean af tangent line is given as d-fis)= f'ls) (at-s) - 2x3 = 39= 21-8